ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

8. Analysis

1991 MS Classification: 33, 44, 45 Kumbhat, R. K., JNV University, Jodhpur, India Some Special Functions of Matrix Arguments The aim of the present paper is to give the integral representation for generalized hypergeometric Functions F_{N}, F_{S} and F_{T} of three arguments when the matrices are real symmetric positive definite. Various theorems connecting these functions with hypergeometric functions of matrix arguments have also been obtained.